OpenStudy (anonymous):

c: Algebra 1 Questions

OpenStudy (anonymous):

Here's my first question. ^_^

OpenStudy (texaschic101):

3r + 2g 3(5) + 2(6) = 15 + 12 = 27

OpenStudy (anonymous):

part a. 3r+2g=p part b. 3*5+2*6 15+12=27

OpenStudy (texaschic101):

great minds think alike :)

OpenStudy (anonymous):

i guess so

OpenStudy (anonymous):

O: I see. Quite helpful! ^-^

OpenStudy (anonymous):

part a. 35+50x<1325 part b. 35+50x<1325 -35 50x<1290 divide by 50 x<25.8 so 25 monthes

OpenStudy (texaschic101):

35 + 50m <= 1325 35 + 50m = 1325 50m = 1325 - 35 50m = 1290 m = 1290/50 m = 25.8 answer would be 25

OpenStudy (anonymous):

I see. For a comedian, you are quite the mathematician! And you're awesome too texaschic. ^-^

OpenStudy (texaschic101):

I am not good with functions...sorry...your gonna have to go with mathman's answer

OpenStudy (anonymous):

thanks thisisausername part a. 900-75=825 part b. 900-75*10=900-750=150 part c. 900-75x=0 -900 -75x=-900 divide by -75 x=12

OpenStudy (texaschic101):

you got to write it as a function....f(x) = 900 - 75x Does that look right ??

OpenStudy (anonymous):

for part c u need a function not for any other part though

OpenStudy (texaschic101):

oh...I thought you needed it for part A. and x would be 1(for 1 month)

OpenStudy (anonymous):

well if you want the extra work

OpenStudy (anonymous):

I see... very good work mathman! ^-^

OpenStudy (anonymous):

(9x)^2=9x*9x=81x^2 squareroot(9x^4)=Squareroot(9)*squareroot(x^4)=3*x^2=3x^2

OpenStudy (anonymous):

Is the top one A, and the bottom one B?

OpenStudy (texaschic101):

yes

OpenStudy (anonymous):

^^ yes

OpenStudy (anonymous):

Ah, thanks much my good friend. I have but one more question. You are both lifesavers. c:

OpenStudy (texaschic101):

t + c <= 20 15t + 8c >= 150

OpenStudy (anonymous):

\[t+c \le 20\]\[15t+8c \ge150\]

OpenStudy (texaschic101):

confirmation is always good :)

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

^-^ Thank you guys so much! You made my day awesome! I hope you guys have a good day too. :D

OpenStudy (texaschic101):

you too :)

OpenStudy (anonymous):

no problem

OpenStudy (texaschic101):

mathman...it shouldn't take you long to get a real high smart score because you are good :)

OpenStudy (anonymous):

um i've been here for about 5-6 monthes min and i have a score of 64